Grade 12Algebra

Find the ratio of the following:
55 paise to ₹ 1.............

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

55 paise to₹ 1
₹1 = 100 paise
Required ratio = 55 / 100 = (11×5) / (20 × 5)
=11 / 20

Thanks